OpenStudy (anonymous):

-1/4w - 1=6 show work please

OpenStudy (saifoo.khan):

\[-\frac 14w = 6 + 1\] \[-\frac 14w = 7\] Cross multiply, \[-w = 7 \times 4\] \[-w = 28\] \[w = -28\]